Two cars traveling in the same direction leave the same point at the same time. Car 1 is going 65 km/h and Car 2 is going 55 km/h. How far must Car 1 travel before it has a 15 min. lead on Car 2.
Answer by mananth(16949)

since the cars are travelling in same direction Car 1 gains 10 km every hour.
in 1/4 hour it will gain 10/(1/4)= 2.5 km.
..
Let car 1 travel x km at speed of 65 km/h
car2 will travel x-2.5 km in the same time
..
time = distance /speed
..
x/65 = (x-2.5)/55
55x=65(x-2.5)
55x=65x-162.5
add -65x
55x-65x=65x-65x-162.5
-10x=-162.5
divide by -10
x= -162.5/-10
x=16.25 km. the distance car 1 has to travel to be 15 minutes ahead.